OpenStudy (anonymous):

The Oliver Company plans to market a new product. Based on its market studies, Oliver estimates that it can sell up to 5,500 units in 2005. The selling price will be $3 per unit. Variable costs are estimated to be 10% of total revenue. Fixed costs are estimated to be $10,800 for 2005. How many units should the company sell to break even?

OpenStudy (anonymous):

Your figure for the fixed costs is somewhat confusing. I will assume you mean $10,800 Revenue Function: R(x) = 3x Cost Function: C(x) = 0.1[R(x)] + 10800 C(x) = 0.1(3x) + 10800 C(x) = 0.3x + 10800 Break-Even Point: R(x) = C(x): 3x = 0.3x + 10800 3x - 0.3x = 10800 2.7x =10800 x = 10800 / 2.7 x = 4000 Break-Even Point = 4,000 units ¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯ Hope this helps